Keith, Michael Circle digits : a self referential story

For a time

I stood pondering on circle sizes. The large computer mainframe quietly proces-

sed all of its assembly code. Inside my entire hope lay for figuring out an elusive expansion. Value : pi. Decimals

expected soon. I nervously entered a format procedure. The mainframe processed the request. Error. I, again entering it, carefully

retyped. This iteration gave zero error printouts in all - success. Intently I waited. Soon, roused by thoughts within me, appeared narrative mnemonics relating digits to verbiage ! The idea appeared to exist but only in abbreviated fashion - little phrases typically. Pressing on I then resolved, deciding firmly about a sum of decimals to use - likely around four hundred, presuming the computer code soon halted ! Pondering these ideas, words appealed to me. But a problem of zeros did exist. Pondering more, solution subsequently appeared. Zero suggests a punctuation element. Very novel ! My thoughts were culminated. No periods, I concluded. All residual marks of punctuation = zeros. First digit expansion answer then came before me. On examining some problems unhappily arose. That imbecillic bug ! The printout I processed showed four nine as foremost decimals. Manifestly troubling. Totally every number looked wrong. Repairing the bug took much effort. A pi mnemonic with letters truly seemed good. Counting of all the letters probably should suffice. Reaching for a record would be helpful. Consequently, I continued, expecting a good final answer from computer. First number slowly displayed on the flat screen -3. Good. Trailing digits apparently were right also. Now my memory scheme must be probably implementable. The technique was chosen, elegant in scheme : by self reference a tale mnemonically helpful was ensured. An able title suddenly existed - ”Circle Digits”. Taking pen I began. Words emanated uneasily. I desired more synonyms. Speedily I found my (alongside

me) Thesaurus. Rogets is probably an essential in doing this, instantly I

decided. I wrote and erased more. The Rogets clearly assisted immen-

sely. My story proceeded (how lovely !) faultlessly. The end, above

all, would soon joyfully overtake. So, this memory helper

story is incontestably complete. Soon I will locate

publisher. There a narrative will I trust imme-

diately appear producing fame. The end.